OpenStudy (anonymous):

solve for the variable. reduce the answer to lowest terms p/-25=-650

OpenStudy (anonymous):

-25*(p/-25) = -650*-25 p = 16250

OpenStudy (anonymous):

Multiply by (-25) both the sides, you will get the value of p ie p = -650*(-25) = 650*25 multiply it and get the answer..

OpenStudy (anonymous):

Your goal is to isolate the variable. You should starting trying the answers yourself and then post if you are not confident and people will tell you if you get them right or wrong. This way you can start to learn.

OpenStudy (anonymous):

start**